# Data-Driven Predictive Control for Robust Exoskeleton Locomotion

### v1

@@ V-B Hardware Results
Different User Settings: We conducted two sets of experiments to test the framework with different payloads and kinematics as this platform is designed to be used by different users. The data collection part for the hardware experiment is similar to that of the simulation setup for tracking performance comparison with gaits at different step lengths. The first set of experiments is conducted with the exoskeleton carrying 202020 (kg) of payload with the link lengths of a subject of 1.741.741.74 (m). DDPC planner is able to regulate the CoP position to a more centered position as depicted in Fig. 8a and 8b. A different set of hyperparameters with $T=800$, $T_{\textrm{ini}}=20$, $N=100$, and $\delta_{t}=0.015$ is used. The second experiment is conducted with a user of height 1.631.631.63 (m) and 525252 (kg) with the same hyperparameter to construct the Hankel Matrix as in tracking performance case. The gait tiles for the experiment with different payloads are shown in Fig. 8c and 8d.
@@ VI CONCLUSIONS
This paper successfully demonstrated the DDPC framework’s application on lower-body exoskeleton, both in simulations and on hardware. Through detailed simulation analyses, the DDPC framework proved its effectiveness in stabilizing the system, surpassing traditional physics-based template models such as LIP, particularly at increased desired speeds. The framework’s robustness was also validated on hardware, showcasing its ability to accommodate model discrepancies beyond the initial model used for data collection. Furthermore, we introduced a time-varying perturbation in the simulation while updating the transition matrix online. Although further research is necessary to refine the online update process of the Hankel matrices systematically, these results underscore the framework’s capacity to adapt to changing environments. However, it is crucial to note that the DDPC planner’s performance is intrinsically linked to the trajectory used to construct the Hankel matrices. It can only enhance and build upon the capabilities of the nominal controller used for data collection. Moreover, considering the inherent limitations in the actuation of the CoM horizontal position, future studies will explore how to extend the proposed framework to incorporate foot placement and step timing planning to enhance stability and performance.

### v2

@@ V-B Hardware Results
Different User Settings: We conducted two sets of experiments to test the framework with different payloads and kinematics as this platform is designed to be used by different users. The data collection part for the hardware experiment is similar to that of the simulation setup for tracking performance comparison with gaits at different step lengths. The first set of experiments is conducted with the exoskeleton carrying $20$ kg of payload with the link lengths of a subject of $1.74$ m. DDPC planner is able to regulate the CoP position to a more centered position as depicted in Fig. 8a and 8b. A different set of hyperparameters with $T=800$, $T_{\textrm{ini}}=20$, $N=100$, and $\delta_{t}=0.015$ is used. The second experiment is conducted with a user of height $1.63$ m and $52$ kg with the same hyperparameter to construct the Hankel Matrix as in tracking performance case. The gait tiles for the experiment with different payloads are shown in Fig. 8c and 8d.
@@ VI CONCLUSIONS
This paper successfully demonstrated the DDPC framework’s application on lower-body exoskeleton, both in simulations and on hardware. Through detailed simulation analyses, the DDPC framework proved its effectiveness in stabilizing the system, surpassing traditional physics-based template models such as LIP, particularly at increased desired speeds. The framework’s robustness was also validated on hardware, showcasing its ability to accommodate model discrepancies beyond the initial model used for data collection. Furthermore, we introduced a time-varying perturbation in the simulation while updating the transition matrix online. Although further research is necessary to refine the online update process of the Hankel matrices systematically, these results underscore the framework’s capacity to adapt to changing environments. However, it is crucial to note that the DDPC planner’s performance is intrinsically linked to the trajectory used to construct the Hankel matrices. It can only enhance and build upon the capabilities of the nominal controller used for data collection. Moreover, considering the inherent limitations in the actuation of the CoM horizontal position, future studies will explore how to extend the proposed framework to incorporate foot placement and step timing planning to enhance stability and performance. Additionally, further investigation into incorporating more information on user movement would be beneficial.